WHAT MOLECULAR LINES CAN TELL ABOUT EARLY STAGES OF MASSIVE STARS

TATIANA VASYUNINA, ERIC HERBST, Ohio State University, 191 W. Woodruff Ave.,43210, Columbus, OH, USA; HENDRIK LINZ, THOMAS HENNING, HENRIK BEUTHER, Max Planck Institute for Astronomy (MPIA), Königstuhl 17, D-69117 Heidelberg, Germany; IGOR ZINCHENKO, Institute of Applied Physics of the Russian Academy of Sciences, Ulyanova 46, 603950 Nizhny Novgorod, Russia; MAXIM VORONKOV, Australia Telescope National Facility, CSIRO Astronomy and Space Science, PO Box 76, Epping, NSW 1710, Australia.

Here we present our molecular line investigation of the southern infrared dark clouds (IRDCs). We performed observations of 13 molecular species using the 22-m Mopra radio telescope. In our survey we included in general species which are good tracers of cold and dense gas. Some of them trace the more quiescent gas (e.g. HNC, N_{2}H^{+}), while others are sensitive to more dynamical processes (HCN, HCO^{+}). We detect SiO emission in some clouds and complicated shapes of the HCO^{+} emission line profile in all IRDCs, which indicates infall and outflow motions and the beginning of starformation activity, at least in some parts of the IRDCs. Using H_{2} column densities from our previous investigation, we estimated molecular abundances for all species. We uncovered a tendency for IRDCs to have molecular abundances similar to those in low-mass pre-stellar cores. This similarity may indicate similar chemical composition at the earliest stages of low- and high-mass stars and their close evolutionary status. However, for the N_{2}H^{+} and HCO^{+} species, there is a tendency for higher mean abundances in IRDCs compared with low- mass cores. To find a reason of such behavior, we use chemical models including gas-phase reactions and accretion and desorption onto/from grains. Chemical modeling allows us to study chemical evolution at the early stages of massive stars in more detail and reconstruct the physical conditions and evolutionary status of IRDCs.